OpenStudy (anonymous):

Write the equation of the axis of symmetry,and find the coordinates of the vertex of the graph of each function.Identify the vertex as a maximum or minimum. Then graph the function Y=-3X^2-6X+4

OpenStudy (anonymous):

y=-3(x^2+2x)+4 =-3(x^2+2x+1-1)+4 =-3(x+1)^2+3+4 y-7=-3(x+1)^2 axis of symmetry is x+1=0 vertex is (-1,7) it is a downward parabola with maxima at (-1,7) and maximum value of y=7